# Efficient Bayesian Computation Using Plug-and-Play Priors for Poisson Inverse Problems

### v1

@@ B.2 Additional results for PnP-MLA using Prox-DRUNet with $\epsilon=20$
Here, additional results to Section 4.3 are presented. We use a detail of the starfish image shown in Figure 15 to illustrate the effects of the chosen step size. Results for the PnP-MLA using $\epsilon=20$ are summarized in Table 8, yielding similar results to PPnP-ULA in Table 5 in terms of best metric and step size, both quantitatively and qualitatively (compare Figures 8.2 and 16, with the best results in Figure 16(b)). With increasing step size the algorithm shows instability by introducing artefacts that only appear in the starfish image (not in any other image in set3c - not shown here), see Figure 16(d) and 16(e). We find that by choosing $\epsilon=25$ for the neural network denoiser (these results are described in the main text for Section 4.3), and thus increasing the denoising strength slightly, all artefacts can be mitigated.

### v2

@@ B.2 Additional results for PnP-MLA using Prox-DRUNet with $\epsilon=20$
Here, additional results to Section 4.3 are presented. We use a detail of the starfish image shown in Figure 17 to illustrate the effects of the chosen step size. Results for the PnP-MLA using $\epsilon=20$ are summarized in Table 8, yielding similar results to PPnP-ULA in Table 5 in terms of best metric and step size, both quantitatively and qualitatively (compare Figures 8.2 and 18, with the best results in Figure 18(b)). With increasing step size the algorithm shows instability by introducing artefacts that only appear in the starfish image (not in any other image in set3c - not shown here), see Figure 18(d) and 18(e). We find that by choosing $\epsilon=25$ for the neural network denoiser (these results are described in the main text for Section 4.3), and thus increasing the denoising strength slightly, all artefacts can be mitigated.